Two charges Q1 (= +6 μC) and Q2 (= –2 μC) are brought from infinity to positions on the x-axis of x = –4 cm and x = +4 cm, respectively. How much work was done in bringing the charges together?
Answer #1

Somo - Work done(w) =kQ, Q₂ ४ 2 Whese r= separation distance = ucmtucm = 8cm = 0.08m k=gX10 Norg (2 iwa (9X10 N. m ) (6X15%()
Answer #2

SOLUTION : 


Q1 = 6 µ C = 6 * 10^(-6) C

Q2 = - 2 µ C = - 2 * 10^(-6) C 


Distance between Q1 and Q2 , r = 4 - (-4) = 8 cm = 0.08 m^2 


W = k Q1 Q2 / r 


Where k = 9.0 * 10^9 Nm^2 / C^2 


So,


= 9.0 * 10^9 * (6 * 10^(-6)) * (- 2 * 10^(-6) / 0.08  (N-m)

= - 1.35 N-m

= - 1.35 J (ANSWER)

answered by: Tulsiram Garg